OpenStudy (anonymous):

Let f(x)=x^2-(1/sqrtx). What is the x-coordinate of the point of inflection?

OpenStudy (anonymous):

\[f(x)=x^2-\frac{1}{\sqrt{x}}\]

OpenStudy (anonymous):

you need second derivative for this one

OpenStudy (anonymous):

Should I rewrite it first?

OpenStudy (anonymous):

you can if that helps. write it as \[x^2-x^{-\frac{1}{2}}\]

OpenStudy (anonymous):

then \[2x+\frac{1}{2}x^{-\frac{3}{2}}\] for the first one

OpenStudy (anonymous):

okay I think I know how to work it now

OpenStudy (anonymous):

ok good, but you have to get it out of exponential notation at the end so you can set it equal zero and solve

OpenStudy (anonymous):

okay well lets see what I get.

OpenStudy (anonymous):

so I'm stuck at \[x^(-5/2)=8/3\]

OpenStudy (anonymous):

Where did I go wrong?